OpenStudy (anonymous):

Solve for x. x -3/8 = -5/8 x = 1/4 x = -1/4 x = -1 x = 1

OpenStudy (anonymous):

Just bring -3/8 to the otherside. What's x = -5/8 + 3 /8?

OpenStudy (anonymous):

x = -2/8, which can be reduced to -1/4.